Swiss Supplies showed the following selected adjusted balances at its December 31, 2019, year-end: Accounts Receivable Dec. 3Part B Required: 2. Prepare the December 31, 2020, adjusting entry to estimate bad debts assuming that uncollectible accounts3. Show how accounts receivable will appear on the December 31, 2020, balance sheet. Partial Balance Sheet Assets Current assPart C (independent of Part B) Required: 5. Prepare the December 31, 2020, adjusting entry to estimate bad debts assuming tha6. Show how accounts receivable will appear on the December 31, 2020, balance sheet. Partial Balance Sheet Assets Current ass
Answer #1

1)

Date General Journal Debit Credit
a) Cash (2780000*25%) 695000
Accounts Receivable 2085000
Sales 2780000
Cost of goods sold 1795000
Inventory 1795000
b) Sales return 118000
Accounts receivable 59000
Cash 59000
c) Cash 34000
Accounts receivable 34000
d) Bad debts 36000
Accounts receivable 36000
e) Cash 1740000
Accounts receivable 1740000

2)

Date General Journal Debit Credit
Dec-31 Bad debt expense 20260
Allowance for doubtful debts 20260

Net credit sales = 2085000-59000 = 2026000

Allowance for doubtful debts = 2026000*1% = 20260

3)

Partial Balance Sheet
Assets
        Current Assets:
             Accounts Receivable 606000
             Less: Allowance for doubtful debts 20260 585740

Outstanding accounts receivable = 390000+2085000-59000-34000-36000-1740000 = 606000

4) Bad debt expense = 20400 + 20260 + 36000 = 76660

5) Outstanding receivables = 606000*3% = 18180

Date General Journal Debit Credit
Dec-31 Bad debt expense 18180
Allowance for doubtful debts 18180

6)

Partial Balance Sheet
Assets
        Current Assets:
             Accounts Receivable 606000
             Less: Allowance for doubtful debts 18180 587820

7)

Bad debt expense = 20400 + 18180 +36000 = 74580